OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 6 and contains the point (1, 1).

OpenStudy (anonymous):

for the line with slope m and passing through (x1,y1) use (y-y1)=m(x-x1)

OpenStudy (anonymous):

y = mx + b, m = slope = 6. \[y = 6x + b\] 1 = 6(1) + b \[b = -5\] y = 6x-5

OpenStudy (anonymous):

aliter y-1 =6(x-1) y-1=6x-6 y=6x-6+1 or y=6x-5